Hi, how do you solve x/4-y/4=-1 and x+4y=-9 by substitution?
Answer by JBarnum(2146)

x/4-y/4=-1
I don't care for fractions so multiply by 4
x-y=-4
Lets solve for x
x=y-4
Substitute what x equals into below equation

x+4y=-9
(y-4)+4y=9
Solve for y
5y=13
y=13/5
x=y-4
x=13/5 - 4
x=13/5 - 20/5
x=-7/5